OpenStudy (anonymous):

(medal!!!!) Joanna wants to buy a car. Her parents loan her $5,000 for 5 years at 5% simple interest. How much will Joanna pay in interest?

OpenStudy (anonymous):

find 5% of 5,000 then multiply by five.... need i say more?

OpenStudy (sbuck98):

5,000*5%=250*5=$1250 in interest.

OpenStudy (anonymous):

Thanks more on the way